# Bounds on the realizations of zero-nonzero patterns and sign conditions of polynomials restricted to varieties and applications

### v2

@@ 4.1. Proof of Theorem 1
In our proof of Theorem 1 we will need to use the following generalized form of Bezout’s theorem (see for instance [18, Example 8.4.6, page 148]).
Theorem I (Generalized Bezout).
@@ 4.1.1. Outline of the proof Theorem 1
The main idea (which is very similar to that used in [25], and also in [8]) is to first bound the number of zero-non-zero patterns whose realizations are zero dimensional using Theorem I. The main subtlety here is that the dimension of an algebraic set need not decrease upon intersection with another hypersurface. The problem of bounding the number of zero-nonzero patterns whose realizations have dimension $>0$ then can be reduced to the zero-dimensional case by taking generic hyperplane sections.
@@
The claim is trivially true for $\ell=1$. Suppose it is true up to $\ell-1$. Then using Theorem I we get that
@@ 4.2.1. Multiplicity and singular points of a variety
Given a variety $V$, and a point $x\in V$, the multiplicity of $V$ at $x$, denoted $\mathrm{mult}_{x}(V)$, is defined as the degree of the exceptional divisor of the blowup $\widetilde{V}$ of $V$ at $x$ [18, Section 4.3, page 79]. This definition as equivalent to the definition of the Samuel multiplicity of the local ring $\mathcal{O}_{V,x}$ [18, Example 4.3.1].
@@
Given a variety $V$, and a point $x\in V$, $\mathrm{mult}_{x}(V)=1$ if and only if the local ring $\mathcal{O}_{V,x}$ is regular [18, Example 4.3.5 (d)], and this is equivalent to $V$ being non-singular at $x$ [18, Section B.2.7, Page 429].
@@
The set of points of $V$ which are not non-singular will be denoted by $\mathrm{Sing}(V)$. It is a proper closed subset of $V$ (see [24, Theorem 5.3]). If $V$ is a variety, then $\dim\mathrm{Sing}(V)<\dim V$.
@@ 4.2.2. Real closed extensions and Puiseux series
If $\mathrm{R}^{\prime}$ is a real closed extension of a real closed field $\mathrm{R}$, and $S\subset\mathrm{R}^{k}$ is a semi-algebraic set defined by a first-order formula with coefficients in $\mathrm{R}$, then we will denote by ${\rm Ext}(S,\mathrm{R}^{\prime})\subset\mathrm{R}^{\prime k}$ the semi-algebraic subset of $\mathrm{R}^{\prime k}$ defined by the same formula. It is well known that ${\rm Ext}(S,\mathrm{R}^{\prime})$ does not depend on the choice of the formula defining $S$ [7, Proposition 2.87].
@@ 4.2.3. Outline of the proof of Theorem 2
We first observe that it suffices to consider the case when $V$ is irreducible of degree $D$ and dimension $p$, satisfying that $\dim_{x}V(\mathrm{R})=p$ for all $x\in V(\mathrm{R})$. We introduce an infinitesimal $\gamma_{0}$. Using the conic structure theorem for semi-algebraic sets [7, Proposition 5.49] we are able to restrict attention to realizations of sign conditions which are contained in a ball of radius $(1/\gamma_{0})^{1/2}$. This necessitates introduction of a new polynomial denoted by $P_{0}$ in the proof. We then define infinitesimal perturbations (using $s+2$ additional infinitesimals $\gamma_{1},\ldots,\gamma_{s},{\varepsilon},\delta$) of the polynomials in $\mathcal{P}$ and prove that it suffices to bound the number of semi-algebraically connected components of the realizations on $\widetilde{S}:={\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}})$ (the extension of $V(\mathrm{R})$ to the field $\widetilde{\mathrm{R}}$ obtained from $\mathrm{R}$ by adjoining infinitesimal elements) of a certain set of weak sign conditions on this new set of $14s+$ 1 polynomials (Lemma 4.4). We then show that in order to bound these, it suffices to bound the number of semi-algebraically connected components of all the non-empty real algebraic sets defined by the perturbed set of $14s+$ 1 polynomials which are contained in $\widetilde{S}$ (Lemma 4.7).
@@ 4.2.4. Proof of Theorem 2
Let $x\in C$. Then, $x\in{\mathcal{R}}(\widetilde{\phi}_{\sigma},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))$, and let $\widetilde{C}_{\sigma}$, be the semi-algebraically connected component of ${\mathcal{R}}(\widetilde{\phi}_{\sigma},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))$ containing $x$. Then $\widetilde{C}_{\sigma}\cap V(\mathrm{R})=C\cap V(\mathrm{R})$ (see [7, Proposition 13.7]). ∎
@@
Let $\mathcal{Q}^{\prime}$ be a maximal subset of $\mathcal{Q}$ with the property that ${\rm Z}(\mathcal{Q}^{\prime},\mathrm{R}^{N})\cap C\neq\emptyset$. Then $\mathcal{Q}^{\prime}$ satisfies the required property in the lemma (see [7, Proposition 13.1]). ∎
@@
Let $\widetilde{S}$ be the intersection of ${\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})$ with the set defined by $-\zeta\leq Q\leq\zeta$. Let $C\in\mathrm{Cc}({\rm Z}(Q,S))$ and $x\in C$. Then, it follows from the hypothesis that $\dim_{x}{\rm Z}(Q,C)<\dim_{x}S$. Hence, for all $r>0$, denoting (here and elsewhere) by $B_{N}(x,r)$ the open Euclidean ball centered at $x$ of radius $r$, ${\rm Z}(Q,C)\cap B_{N}(x,r)$ is non-empty and properly contained in $S\cap B_{N}(x,r)$. So $S-{\rm Z}(Q,C)\neq\emptyset$ and $x\in\overline{S-{\rm Z}(Q,C)}$. Applying the semi-algebraic curve selection lemma (see for example [7, Theorem 3.19]), there exists a semi-algebraic curve $\gamma:[0,1]\rightarrow S$, such that $\gamma(0)=x$, and $\gamma((0,1])\subset S-{\rm Z}(Q,C)$. In particular, $Q(\gamma(1))\neq 0$.
Suppose without loss of generality that $Q(\gamma(1))>0$. Using the intermediate value property of the field $\mathrm{R}{\langle}\zeta{\rangle}$ ([7, Proposition 4.3]), there exists $\widetilde{t}\in\mathrm{R}{\langle}\zeta{\rangle},\widetilde{t}>0$, such that $Q({\rm Ext}(\gamma,\mathrm{R}{\langle}\zeta{\rangle})(\widetilde{t}))=\zeta$, and let $\widetilde{x}={\rm Ext}(\gamma,\mathrm{R}{\langle}\zeta{\rangle})(\widetilde{t})$. Then $\widetilde{x}\in{\rm Z}(Q-\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle}))$. Let $\widetilde{C}\in\mathrm{Cc}({\rm Z}(Q-\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})))$ with $\widetilde{x}\in\widetilde{C}$. Then, $\lim_{\zeta}\widetilde{C}\subset C$. Thus we have proved that for each $C\in\mathrm{Cc}({\rm Z}(Q,S))$, there exists $\widetilde{C}\in\mathrm{Cc}({\rm Z}(Q-\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})))\cup\mathrm{Cc}({\rm Z}(Q+\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})))$, with $\lim_{\zeta}\widetilde{C}\subset C$. The lemma follows. ∎
@@
It follows from the Noether normalization theorem (see for example [15, Theorem 2.2.9]) that for a generic linear projection map $P:\mathrm{C}^{N}\rightarrow\mathrm{C}^{p}$ of rank $p$, $V_{\ell}:=P^{-1}(\ell)\cap V$ is (non-empty and) finite for all $\ell\in L=\mathrm{Im}(P)\subset\mathrm{C}^{N}$.
@@
It follows from Noether normalization theorem (see for example [15, Theorem 2.2.9]) that for $\pi$ generic, $\pi(V)$ is a hypersurface in $\mathrm{C}^{p+1}$.
@@
using the basic inequality on the Betti numners of real algebraic sets due to Oleĭnik and Petrovskiĭ [34] (see for example [7, Proposition 7.28]). The lemma follows. ∎
@@ 4.3.2. Proof of Theorem 5
Let $K=\bigcup_{Q\in\mathcal{Q}}{\rm Z}(Q^{h},\overline{V}(\mathrm{R}))$. Since $\overline{V}(\mathrm{R})$ is a compact (real) manifold (orientable over $\mathbb{Z}/2\mathbb{Z}$, and $K$ is a closed and bounded semi-algebraic subset, we obtain from Lefschetz duality theorem (see for example [43, page 297, Theorem 18]),  [^4] By a standard application of the Tarski-Seidenberg transfer principle we can assume that it holds over a general real closed field $\mathrm{R}$. that for each $i$, $\mbox{\rm H}_{i}(\overline{V}(\mathrm{R})-K)\cong\mbox{\rm H}^{\dim\overline{V}-i}(\overline{V}(\mathrm{R}),K)$. Moreover, using the cohomology long exact sequence one has
@@
If $X$ is a compact space (or a regular complex) equipped with an involution map $c:X\rightarrow S$, and $\mathrm{Fix}(c)\subset X$ denotes the subspace of fixed points of $X$. The Smith exact sequence (see for example [14, page 126]) then implies that
@@
If $\overline{V}\subset\mathbb{P}^{N}_{\mathrm{C}}$ is a non-singular complete intersection variety with $\dim\overline{V}=N-\ell$, and defined as the intersection of non-singular hypersurfaces of degrees $e_{1},\ldots,e_{\ell}$, then $b(\overline{V})$ is given by the following classical formula (see for example [16, §5.7.2-3])
@@
In our case, we can assume that the degree sequence of the polynomials defining $\overline{V}$ is of length $\ell=N-p+q$. Suppose that the degree sequence is $e_{1},\ldots,e_{N-p+q}$. We can assume without loss of generality that $e_{1}\cdot e_{N-p}=D$, and $e_{N-p+1}\leq d,\ldots,e_{N-p+q}\leq d$.
@@
First observe that it follows from the theorem of conic structure at infinity [7, Proposition 5.49], that for all large enough $R>0$,
@@ 4.3.4. Proof of Theorem 6
Now let $V\subset\mathrm{C}^{N}$ be a complete intersection variety of dimension $p$ defined by polynomials $\mathcal{Q}=\{Q_{1},\ldots,Q_{N-p}\}\subset\mathrm{C}[Z_{1},\ldots,Z_{N}]$, and let $\widetilde{\mathcal{Q}}$ be the set of corresponding polynomials in Lemma 4.20.
@@ 4.4.2. Proof of Theorem 7
We already have the necessary ingredients to start the proof without having to proof additional lemmas.
@@
Using the conic structure at infinity for semi-algebraic sets (see for example [7, Proposition 5.49]) choose $R>0,R\in\mathrm{R}$, such that for each $\sigma\in\{0,1,-1\}^{\mathcal{P}}$, ${\mathcal{R}}(\sigma,V(\mathrm{R}))$ is semi-algebraically homeomorphic to ${\mathcal{R}}(\sigma,V(\mathrm{R})\cap B_{N}(0,R))$. We denote $\mathcal{P}_{0}=\mathcal{P}\cup P_{0}$, where $P_{0}=\sum_{i=1}^{N}X_{i}^{2}-R$.
@@
Let $H\in\widetilde{\mathrm{R}}[X_{0},\ldots,X_{N}]$ be a generic homogeneous polynomial of the same degree as $\widetilde{Q}$ and which in non-negative in $\mathrm{R}^{N+1}$. Note that in $\mathbb{P}_{\mathrm{R}}^{N}$ the sign of a homogeneous polynomial of even degree is well defined. It follows from [7, Lemma 16.17] that the semi-algebraic set $\widetilde{S}\subset{\rm Ext}(\overline{V}(R),\widetilde{R}{\langle}\zeta{\rangle}))$ defined as the intersection of ${\rm Ext}(\overline{V}(R),\widetilde{R}{\langle}\zeta{\rangle}))$ with the semi-algebraic set defined by $(1-\zeta)\widetilde{Q}-\zeta H\leq 0$ is semi-algebraically homotopy equivalent to ${\rm Ext}(S,\widetilde{\mathrm{R}}{\langle}\zeta{\rangle})$.
@@
4.5. Proofs of Theorems 8 and 9
Proof of Theorem 8.
We only consider the affine case. The projective case can be reduced to the affine case using by taking affine covers. For each $n>0$, let
$\mathcal{Q}_{n}\subset k[\mathbf{X}^{(1)},\ldots,\mathbf{X}^{(n)}],$
be defined by
$\mathcal{Q}_{n}=\bigcup_{1\leq i,j\leq n}\bigcup_{P\in\mathcal{P}_{n}}P(\mathbf{X}^{(i)},\mathbf{X}^{(j)}),$
where each $\mathbf{X}^{(i)}$ is a block of $N_{n}$-variables.
Then, then it is easy to verify that the of number distinct $(Z_{n},F_{n})$-graphs is bounded by $\mathrm{card}(\Pi(\mathcal{Q}_{n},Z_{n}^{n}))$.
Now, $\deg(Z_{n}^{n})=\deg(Z_{n})^{n}\leq c_{3}^{n}n^{n}$, $\dim Z_{n}^{n}\leq np$, $\mathrm{card}(\mathcal{Q}_{n})=c_{1}n^{2}$, $\deg(Q)\leq c_{2}n$ for $Q\in\mathcal{Q}_{n}$.
Applying Theorem 1 we obtain
$\mathrm{card}(\Pi(\mathcal{Q}_{n},Z_{n}^{n}))$$\leq$ $c_{3}^{n}n^{n}\sum_{i=0}^{np}\binom{c_{1}n^{2}}{i}(c_{2}n)^{pn-i}$
$\leq$ $c_{3}^{n}n^{n}\sum_{i=0}^{np}\left(\frac{ec_{1}n^{2}}{i}\right)^{i}(c_{2}n)^{pn-i}$
$=$ $c_{3}^{n}\cdot n^{n}\cdot\sum_{i=0}^{np}(ec_{1})^{i}\cdot(c_{2})^{pn-i}\cdot n^{pn}\left(\frac{n}{i}\right)^{i}$
$\leq$ $n^{(1+o(1))(p+1)n}.$
Proof of Theorem 9.
The proof is very similar to the proof of Theorem 8 and we omit the details. In the last step, we use Theorem 6 instead of Theorem 1 to bound the number of realizable sign conditions. The bound that we obtain is slightly worse (in the exponent) than the bound in Theorem 8 reflecting the fact that the bound on the realizable sign conditions in Theorem 6 is a bit worse than the one in Theorem 1. ∎

### v3

@@ 4.1. Proof of Theorem 1
In our proof of Theorem 1 we will need to use the following generalized form of Bezout’s theorem (see for instance [15, Example 8.4.6, page 148]).
Theorem H (Generalized Bezout).
@@ 4.1.1. Outline of the proof Theorem 1
The main idea (which is very similar to that used in [21], and also in [6]) is to first bound the number of zero-non-zero patterns whose realizations are zero dimensional using Theorem H. The main subtlety here is that the dimension of an algebraic set need not decrease upon intersection with another hypersurface. The problem of bounding the number of zero-nonzero patterns whose realizations have dimension $>0$ then can be reduced to the zero-dimensional case by taking generic hyperplane sections.
@@
The claim is trivially true for $\ell=1$. Suppose it is true up to $\ell-1$. Then using Theorem H we get that
@@ 4.2.1. Multiplicity and singular points of a variety
Given a variety $V$, and a point $x\in V$, the multiplicity of $V$ at $x$, denoted $\mathrm{mult}_{x}(V)$, is defined is the degree of the exceptional divisor of the blowup $\widetilde{V}$ of $V$ at $x$ [15, Section 4.3, page 79]. This definition is equivalent to the definition of the Samuel multiplicity of the local ring $\mathcal{O}_{V,x}$ [15, Example 4.3.1].
@@
Given a variety $V$, and a point $x\in V$, $\mathrm{mult}_{x}(V)=1$ if and only if the local ring $\mathcal{O}_{V,x}$ is regular [15, Example 4.3.5 (d)], and this is equivalent to $V$ being non-singular at $x$ [15, Section B.2.7, Page 429].
@@
The set of points of $V$ which are not non-singular will be denoted by $\mathrm{Sing}(V)$. It is a proper closed subset of $V$ (see [20, Theorem 5.3]). If $V$ is a variety, then $\dim\mathrm{Sing}(V)<\dim V$.
@@ 4.2.2. Real closed extensions and Puiseux series
If $\mathrm{R}^{\prime}$ is a real closed extension of a real closed field $\mathrm{R}$, and $S\subset\mathrm{R}^{k}$ is a semi-algebraic set defined by a first-order formula with coefficients in $\mathrm{R}$, then we will denote by ${\rm Ext}(S,\mathrm{R}^{\prime})\subset\mathrm{R}^{\prime k}$ the semi-algebraic subset of $\mathrm{R}^{\prime k}$ defined by the same formula. It is well known that ${\rm Ext}(S,\mathrm{R}^{\prime})$ does not depend on the choice of the formula defining $S$ [5, Proposition 2.87].
@@ 4.2.3. Outline of the proof of Theorem 2
We first observe that it suffices to consider the case when $V$ is irreducible of degree $D$ and dimension $p$, satisfying that $\dim_{x}V(\mathrm{R})=p$ for all $x\in V(\mathrm{R})$. We introduce an infinitesimal $\gamma_{0}$. Using the conic structure theorem for semi-algebraic sets [5, Proposition 5.49] we are able to restrict attention to realizations of sign conditions which are contained in a ball of radius $(1/\gamma_{0})^{1/2}$. This necessitates introduction of a new polynomial denoted by $P_{0}$ in the proof. We then define infinitesimal perturbations (using $s+2$ additional infinitesimals $\gamma_{1},\ldots,\gamma_{s},{\varepsilon},\delta$) of the polynomials in $\mathcal{P}$ and prove that it suffices to bound the number of semi-algebraically connected components of the realizations on $\widetilde{S}:={\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}})$ (the extension of $V(\mathrm{R})$ to the field $\widetilde{\mathrm{R}}$ obtained from $\mathrm{R}$ by adjoining infinitesimal elements) of a certain set of weak sign conditions on this new set of $14s+$ 1 polynomials (Lemma 4.4). We then show that in order to bound these, it suffices to bound the number of semi-algebraically connected components of all the non-empty real algebraic sets defined by the perturbed set of $14s+$ 1 polynomials which are contained in $\widetilde{S}$ (Lemma 4.7).
@@ 4.2.4. Proof of Theorem 2
Let $x\in C$. Then, $x\in{\mathcal{R}}(\widetilde{\phi}_{\sigma},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))$, and let $\widetilde{C}_{\sigma}$, be the semi-algebraically connected component of ${\mathcal{R}}(\widetilde{\phi}_{\sigma},{\rm Ext}(V(\mathrm{R}),\widetilde{\mathrm{R}}))$ containing $x$. Then $\widetilde{C}_{\sigma}\cap V(\mathrm{R})=C\cap V(\mathrm{R})$ (see [5, Proposition 13.7]). ∎
@@
Let $\mathcal{Q}^{\prime}$ be a maximal subset of $\mathcal{Q}$ with the property that ${\rm Z}(\mathcal{Q}^{\prime},\mathrm{R}^{N})\cap C\neq\emptyset$. Then $\mathcal{Q}^{\prime}$ satisfies the required property in the lemma (see [5, Proposition 13.1]). ∎
@@
Let $\widetilde{S}$ be the intersection of ${\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})$ with the set defined by $-\zeta\leq Q\leq\zeta$. Let $C\in\mathrm{Cc}({\rm Z}(Q,S))$ and $x\in C$. Then, it follows from the hypothesis that $\dim_{x}{\rm Z}(Q,C)<\dim_{x}S$. Hence, for all $r>0$, denoting (here and elsewhere) by $B_{N}(x,r)$ the open Euclidean ball centered at $x$ of radius $r$, ${\rm Z}(Q,C)\cap B_{N}(x,r)$ is non-empty and properly contained in $S\cap B_{N}(x,r)$. So $S-{\rm Z}(Q,C)\neq\emptyset$ and $x\in\overline{S-{\rm Z}(Q,C)}$. Applying the semi-algebraic curve selection lemma (see for example [5, Theorem 3.19]), there exists a semi-algebraic curve $\gamma:[0,1]\rightarrow S$, such that $\gamma(0)=x$, and $\gamma((0,1])\subset S-{\rm Z}(Q,C)$. In particular, $Q(\gamma(1))\neq 0$.
Suppose without loss of generality that $Q(\gamma(1))>0$. Using the intermediate value property of the field $\mathrm{R}{\langle}\zeta{\rangle}$ ([5, Proposition 4.3]), there exists $\widetilde{t}\in\mathrm{R}{\langle}\zeta{\rangle},\widetilde{t}>0$, such that $Q({\rm Ext}(\gamma,\mathrm{R}{\langle}\zeta{\rangle})(\widetilde{t}))=\zeta$, and let $\widetilde{x}={\rm Ext}(\gamma,\mathrm{R}{\langle}\zeta{\rangle})(\widetilde{t})$. Then $\widetilde{x}\in{\rm Z}(Q-\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle}))$. Let $\widetilde{C}\in\mathrm{Cc}({\rm Z}(Q-\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})))$ with $\widetilde{x}\in\widetilde{C}$. Then, $\lim_{\zeta}\widetilde{C}\subset C$. Thus we have proved that for each $C\in\mathrm{Cc}({\rm Z}(Q,S))$, there exists $\widetilde{C}\in\mathrm{Cc}({\rm Z}(Q-\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})))\cup\mathrm{Cc}({\rm Z}(Q+\zeta,{\rm Ext}(S,\mathrm{R}{\langle}\zeta{\rangle})))$, with $\lim_{\zeta}\widetilde{C}\subset C$. The lemma follows. ∎
@@
It follows from the Noether normalization theorem (see for example [13, Theorem 2.2.9]) that for a generic linear projection map $P:\mathrm{C}^{N}\rightarrow\mathrm{C}^{p}$ of rank $p$, $V_{\ell}:=P^{-1}(\ell)\cap V$ is (non-empty and) finite for all $\ell\in L=\mathrm{Im}(P)\subset\mathrm{C}^{N}$.
@@
It follows from Noether normalization theorem (see for example [13, Theorem 2.2.9]) that for $\pi$ generic, $\pi(V)$ is a hypersurface in $\mathrm{C}^{p+1}$.
@@
using the basic inequality on the Betti numbers of real algebraic sets due to Oleĭnik and Petrovskiĭ [31] (see for example [5, Proposition 7.28]). The lemma follows. ∎
@@ 4.3.2. Proof of Theorem 5
Let $K=\bigcup_{Q\in\mathcal{Q}}{\rm Z}(Q^{h},\overline{V}(\mathrm{R}))$. Since $\overline{V}(\mathrm{R})$ is a compact (real) manifold (orientable over $\mathbb{Z}/2\mathbb{Z}$, and $K$ is a closed and bounded semi-algebraic subset, we obtain from Lefschetz duality theorem (see for example [39, page 297, Theorem 18]),  [^4] By a standard application of the Tarski-Seidenberg transfer principle we can assume that it holds over a general real closed field $\mathrm{R}$. that for each $i$, $\mbox{\rm H}_{i}(\overline{V}(\mathrm{R})-K)\cong\mbox{\rm H}^{\dim\overline{V}-i}(\overline{V}(\mathrm{R}),K)$. Moreover, using the cohomology long exact sequence one has
@@
If $X$ is a compact space (or a regular complex) equipped with an involution map $c:X\rightarrow S$, and $\mathrm{Fix}(c)\subset X$ denotes the subspace of fixed points of $X$. The Smith exact sequence (see for example [12, page 126]) then implies that
@@
If $\overline{V}\subset\mathbb{P}^{N}_{\mathrm{C}}$ is a non-singular complete intersection variety with $\dim\overline{V}=N-\ell$, and defined as the intersection of non-singular hypersurfaces of degrees $e_{1},\ldots,e_{\ell}$, then $b(\overline{V})$ is given by the following classical formula (see for example [14, §5.7.2-3])
@@
In our case, we can assume that the degree sequence of the polynomials defining $\overline{V}$ is of length $\ell=N-p+q$. Suppose that the degree sequence is $e_{1},\ldots,e_{N-p+q}$. We can assume without loss of generality that $e_{1}\cdots e_{N-p}=D$, and $e_{N-p+1}\leq d,\ldots,e_{N-p+q}\leq d$.
@@
First observe that it follows from the theorem of conic structure at infinity [5, Proposition 5.49], that for all large enough $R>0$,
@@ 4.3.4. Proof of Theorem 6
Now let $V\subset\mathrm{C}^{N}$ be a complete intersection variety of dimension $p$ defined by polynomials
$\mathcal{Q}=\{Q_{1},\ldots,Q_{N-p}\}\subset\mathrm{C}[Z_{1},\ldots,Z_{N}],$
and let $\widetilde{\mathcal{Q}}$ be the set of corresponding polynomials in Lemma 4.20.
@@ 4.4.2. Proof of Theorem 7
We already have the necessary ingredients to start the proof without having to prove additional lemmas.
@@
Using the conic structure at infinity for semi-algebraic sets (see for example [5, Proposition 5.49]) choose $R>0,R\in\mathrm{R}$, such that for each $\sigma\in\{0,1,-1\}^{\mathcal{P}}$, ${\mathcal{R}}(\sigma,V(\mathrm{R}))$ is semi-algebraically homeomorphic to ${\mathcal{R}}(\sigma,V(\mathrm{R})\cap B_{N}(0,R))$. We denote $\mathcal{P}_{0}=\mathcal{P}\cup P_{0}$, where $P_{0}=\sum_{i=1}^{N}X_{i}^{2}-R$.
@@
Let $H\in\widetilde{\mathrm{R}}[X_{0},\ldots,X_{N}]$ be a generic homogeneous polynomial of the same degree as $\widetilde{Q}$ and which in non-negative in $\mathrm{R}^{N+1}$. Note that in $\mathbb{P}_{\mathrm{R}}^{N}$ the sign of a homogeneous polynomial of even degree is well defined. It follows from [5, Lemma 16.17] that the semi-algebraic set $\widetilde{S}\subset{\rm Ext}(\overline{V}(R),\widetilde{R}{\langle}\zeta{\rangle}))$ defined as the intersection of ${\rm Ext}(\overline{V}(R),\widetilde{R}{\langle}\zeta{\rangle}))$ with the semi-algebraic set defined by $(1-\zeta)\widetilde{Q}-\zeta H\leq 0$ is semi-algebraically homotopy equivalent to ${\rm Ext}(S,\widetilde{\mathrm{R}}{\langle}\zeta{\rangle})$.